# Reliability, and Global Privacy

### v3

@@ II-D CoGC Framework
    Gradient-Sharing Phase: Client $k$ shares its local model update $\Delta\bm{\Theta}_{k}^{t,I}$ with its neighbors in $\mathcal{V}_{k}=\{m:g_{m,k}\neq 0\}$ (including client $k$ itself).
@@
    Relaying: Client $k$ should listen to its neighbors in $\mathcal{U}_{k}=\{m:g_{k,m}\neq 0\}$ (including client $k$ itself). If client $k$ successfully receives $\Delta\bm{\Theta}_{m}^{t,I}$ from client $m$, it computes the weighted update $g_{k,m}\Delta\bm{\Theta}_{m}^{t,I}$ and aggregates these to form partial sums. If client $k$ successfully hears from all of its $s$ neighbors, the aggregated partial sum is called complete partial sum, otherwise, it is incomplete.
[…]
    Server Aggregation: If the server receives at least KsK-s complete partial sums, it can reconstruct the exact global model in (5) without any distortion, by combining these complete partial sums according to $\bm{C}$.
More details about CoGC can be found in [31].
III-A Method
Consider $K$ clients, where each client $k\in\{1,\cdots,K\}\triangleq[K]$ holds a raw local update $\Delta\bm{\Theta}_{k}^{t,I}$ and a secret key $\bm{N}_{k}^{t}$. To protect data against exposure during transmission, each client locally obscures its raw local model update by applying additive masking, i.e.,
$\bm{Y}_{k}^{t}=\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t}.$ (6)
Then, each client $k$ transmits $\bm{Y}_{k}^{t}$ to its neighbors $m\in\mathcal{V}_{k}$ through unreliable link $\tau_{m,k}^{t}$.
In the meanwhile, client $k$ also acts as a relay, listening to its neighboring client $m\in\mathcal{U}_{k}$. Client $k$ may or may not be able to receive information from client $m$, depending on the realization of link $\tau_{k,m}^{t}$, where
[…]
Let $\mathcal{U}_{k}^{t}$ denote the set of messages that client $k$ can successfully decode at the $t$-th round, i.e., $\mathcal{U}_{k}^{t}=\{m:\tau_{k,m}^{t}=1\}$. Then, client $k$ computes the product $g_{k,m}\bm{Y}_{m}^{t}$ for each $m\in\mathcal{U}_{k}^{t}$, and sums these products to compute the partial sum, i.e.,
@@ III-A Method
If $\mathcal{U}_{k}^{t}=\mathcal{U}_{k}$, meaning that client $k$ successfully decodes the message from all its neighbors, then $\bm{S}_{k}^{t}$ is referred to as a complete partial sum, since the coefficients corresponding to each neighbor’s contribution in $\bm{S}_{k}^{t}$ fully match those in $\bm{G}$.
Then, client $k$ sends $\bm{S}_{k}^{t}$ to the server through link $\tau_{k}^{t}$, where
@@
    Combinator: Let $\bm{c}_{f_{t}}$ denote the $f_{t}$-th row of the combination matrix $\bm{C}$. The combinator $\bm{c}_{f_{t}}$ corresponds to $\mathcal{U}^{t}$ should satisfy
@@
    Server Aggregation: If such $\bm{c}_{f_{t}}$ exists, the server can compute the global model by combining these complete partial sums according to $\bm{c}_{f_{t}}$, i.e.,
@@
Any two: $\bm{\Theta}_{1}+\bm{\Theta}_{2}+\bm{\Theta}_{3}$ $\mathcal{D}_{1}$$\mathcal{D}_{2}$$\mathcal{D}_{3}$ $\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{2}$ $\frac{1}{2}\bm{N}_{1}-\bm{N}_{2}$ $\bm{\Theta}_{2}-\bm{\Theta}_{3}$ $\bm{N}_{2}-\bm{N}_{3}$ $\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{3}$ $\frac{1}{2}\bm{N}_{1}+\bm{N}_{3}$ $\bm{\Theta}_{1}+\bm{N}_{1}$$\bm{\Theta}_{2}+\bm{N}_{2}$$\bm{\Theta}_{3}+\bm{N}_{3}$
Figure 2: An illustrative overview of the proposed secure CoGC framework with simplified notations for clarity.
Core Idea: If the secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ are designed with a special structure such that
$\sum_{k=1}^{K}\bm{N}_{k}^{t}=0,$ (12)
it can be verified that the server aggregation in (11) gives exactly the global model pursued by (5) without introducing any additional ambiguity, as depicted in Fig. 2, i.e.,
@@
During transmission and aggregation, no third party can accurately infer any individual local model. In SecCoGC, if the server receives at least KsK-s complete partial sums, it can fully eliminate the privacy noise and accurately reconstruct the global model. Otherwise, clients proceed to the next round using their latest local models by setting $\bm{\Theta}_{k}^{t+1,0}=\bm{\Theta}_{k}^{t,I}$.
Training terminates when the total number of executed rounds reaches $T$, with the final round resulting in a successful reconstruction of the global model.
III-B Secret Key Generation
Let us define $L$ noise components $\bm{Z}_{l}^{t}\sim\mathcal{N}(\mathbf{0},\bm{I}_{D})$,  [^1] The noise components $\bm{Z}_{l}^{t}$ are not necessarily centered at 0. Here, we assume zero-mean for the convenience of theoretical derivations. where $l\in[L]$ and $\bm{I}_{D}$ is the identity matrix of size $D\times D$. On each client $k$, the secret key is constructed as a combination of these noise components, i.e., $\bm{N}_{k}^{t}=\sum_{l=1}^{L}\alpha_{k,l}\bm{Z}_{l}^{t}$. Let $\bm{A}$ (of size $K\times L$) denote the generator matrix comprising all $\alpha_{k,l}$. To ensure non-zero secret keys, any row in $\bm{A}$ should contain at least one non-zero entry, i.e., $\forall k$, $\exists l:\alpha_{k,l}\neq 0$. In matrix form, the secret keys on all clients are generated by
@@ III-B Secret Key Generation
Next, we begin by reformulating the problem of secure aggregation to accommodate the real field, and then specify the corresponding constraints on the generator matrix $\bm{A}$.
Condition III.1 (Secure Aggregation in $\mathbb{R}$).
First, the secure aggregation scheme must ensure the correct computation of the global model, as specified in (12). This requirement further implies that the sum of the elements in each column of the generator matrix $\bm{A}$ must be zero, i.e.,
@@
Second, any linear combination involving fewer than KK(an adjustable threshold) secret keys must be non-trivial, i.e., the zero vector must be excluded from the span of any such subset so that no combination of fewer than $K$ secret keys cancels out to zero. This requires
@@
This condition guarantees that partial access to fewer than $K$ clients’ information does not undermine security. Note that K1K-1 is the largest possible rank of $\bm{A}$ to satisfy the correctness condition. The rationale behind this threshold $K$ is further discussed below.
Remark III.1 (Necessity of the Security Condition).
The security condition in (16) can be overly strict, as the GC decoding only relies on certain linear combinations of partial sums, while other combinations with incorrect coefficients cannot reveal meaningful information about the underlying model. However, under alternative GC decoding schemes, such as the one proposed in [31], condition (16) remains critical, as it offers privacy guarantees to each individual local model that hold regardless of the employed decoding strategy.
We note that the real-field addition operation cannot give perfect privacy, as the masked version $\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t}$ cannot be fully independent from the original message $\Delta\bm{\Theta}_{k}^{t,I}$. So for any distributions of $\Delta\bm{\Theta}_{k}^{t,I}$ and $\bm{N}_{k}^{t}$, the mutual information is non-zero, i.e., $\forall P_{\Delta\bm{\Theta}_{k}^{t}},P_{\bm{N}_{k}^{t}}:\mathrm{I}(\Delta\bm{\Theta}_{k}^{t,I};\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t})>0$. This property differs fundamentally from that of finite fields, where underlie most existing secure aggregation schemes. For this reason, we adopt $\mu$-local mutual information privacy ($\mu$-LMIP)[5] to measure the privacy level provided by a real-field secure aggregation scheme. The $\mu$-LMIP aligns well with the core objectives of secure aggregation.
Definition III.1 ($\mu$-LMIP).
For a distribution space $P_{\bm{\Theta}}$, a randomized mechanism $\mathcal{M}$ satisfies $\mu$-LMIP iff
$\mathrm{I}(\bm{\Theta};\mathcal{M}(\bm{\Theta}))\leq\mu\;\;\mathrm{bits}.$ (17)
Here, the mechanism $\mathcal{M}$ represents each procedure in SecCoGC, as will be detailed in Section V.
Next, we establish the necessary constraints such that $\bm{A}$ satisfying Condition III.1 exists. The matrix $\bm{A}$ contains $KL$ entries, and (15) imposes $L$ linear constraints. Thus, if $KL\geq L$, which holds for all $K$ and $L$, the system has non-zero solution. Additionally, (16) introduces an extra constraint, requiring $L\geq K-1$. These lead us to the following lemma.
Lemma III.1 (Existence Condition).
A generator matrix $\bm{A}\in\mathbb{R}^{K\times L}$ that gives distinct secret keys satisfying Condition III.1 exists iff
@@
Construction III.1.
Condition III.1 suggests a straightforward method for constructing the generator matrix $\bm{A}$, as outlined below.
@@
The first K1K-1 secret keys are independently generated with full freedom, while the last key is fully dependent on them. Following this, the nature of random generation in $\mathbb{R}$ will ensure the rank condition in (16) with probability $1$.
Example III.1.
Suppose there are $K=5$ clients participating in the training, and $L=K$ independent noise components $\bm{Z}_{l}^{t}\sim\mathcal{N}(\mathbf{0},\bm{I}_{D})$ are used to generate the secret keys. Following Construction III.1, an example of $\bm{A}$ is given below.
$\bm{A}=\begin{bmatrix}1.41&-0.80&0.21&0.72&-0.08\\ 0.29&0.69&-1.16&2.58&-1.93\\ 0.19&0.83&-1.14&-0.66&-0.43\\ 1.58&-0.24&0.10&0.18&-1.79\\ -3.47&-0.48&1.99&-2.82&4.24\\ \end{bmatrix}$. (19)
The secret keys are further generated according to (14). It can be verified that the secret keys satisfy Condition III.1.
III-C Merits and Drawbacks
By inheriting the CoGC framework, SecCoGC maintains computational efficiency and ensures optimality in heterogeneous networks. Moreover, it provides global privacy across all levels of the learning protocol. Specifically, SecCoGC protects against: (i) peer inference attacks, (ii) relay interception, and (iii) server-side inference attacks. Importantly, SecCoGC enables the exact recovery of the global model under unreliable communication, which ensures the model’s utility.
However, Construction III.1 often leads to variations in secret key power. On one hand, signals with very low or high magnitudes are susceptible to distortion or clipping by the typically nonlinear power amplifier, resulting in inaccurate and noisy aggregation. On the other hand, the uneven randomness introduced into local models can lead to unequal levels of privacy preservation across clients. Since learning weights are determined by default (training data), unrelated to factors, e.g., communication and privacy, such disparities in the power of secret keys are unjustified and could raise fairness concerns. The following example concretely illustrates this issue.
Example III.2 (Fairness Issue).
Following Construction III.1, the last key is generated by $\bm{N}_{5}^{t}=-\sum_{k=1}^{4}\bm{N}_{k}^{t}$, so the power of $\bm{N}_{5}^{t}$ can be accumulated significantly. Following Example III.1, we have
@@ III-C Merits and Drawbacks
The last client receives privacy noise with power 3163-16 times higher than others, resulting in disproportionately stronger privacy protection.
IV Toward Fair and Secure Aggregation: Fair-SecCoGC
The condition for fair privacy preservation requires that all secret keys have identical power, i.e., $\forall k\in[K]:\bm{N}_{m}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$. This can be alternatively formulated as follows.
Condition IV.1 (Fairness Condition).
@@ IV Toward Fair and Secure Aggregation: Fair-SecCoGC
The fairness condition in (20) imposes $K$ additional constraints. The solution of $\bm{A}$ satisfying (15) and (20) is non-trivial, if $KL\geq K+L$, which requires $K\geq 2$, $L\geq 2$. Combined with Lemma III.1, we have the following lemma.
Lemma IV.1 (Existence Condition of Fair $\bm{A}$).
A generator matrix $\bm{A}\in\mathbb{R}^{K\times L}$ that gives distinct secret keys and satisfies both the fairness condition (Condition IV.1) and the secure condition (Condition III.1) iff
@@
Furthermore, If $\bm{A}$ is of cyclic structure, the density of the non-zero elements per row is bounded by the following lemma.
Lemma IV.2 (Lower Bound on Density).
If $\bm{A}$ is of cyclic structure, the number of nonzero entries per row should satisfy
@@
Next, we proceed to key construction. To achieve fair secret key generation, the design must simultaneously satisfy the column-wise linear constraints required for security (Condition III.1), and the row-wise $\ell_{2}$-norm constraints required for fairness (Condition IV.1). This dual requirement poses a challenge to code construction, as it necessitates the joint design of all rows and columns, rather than treating them separately. Below, two construction methods are presented.
Construction IV.1 (General Construction).
@@
The (23) can be solved by e.g., the substitution method, the Gröbner Basis Method, etc. However, solving these equations is computationally complex in large systems. To overcome this limitation, we propose a simplified construction of the generator matrix with a cyclic structure, which produces fair and secure secret keys without solving any equations. The core idea is to set the off-diagonal elements identical first, and then compute the diagonal so that the $\ell_{2}$-norm across rows is enforced to be equal. In this case, adjustments of rows do not affect their linear relationships across columns. Under both Construction IV.1 and IV.2, the secret keys are correlated.
Construction IV.2 (Simplified Construction).
Let $L=K$ and each row contains $\gamma\geq 1$ (Lemma IV.2) off-diagonal non-zeros. The following construction of $\bm{A}$ yields fair secret keys.
@@
Example IV.1.
With the same setting in Example III.1, set $\gamma=2$, and $\lambda^{2}=6$, Construction IV.2 gives the following generator matrix,

### v4

@@ III The Proposed Method: Secure Cooperative Gradient Coding (SecCoGC)
           Any two: $\bm{\Theta}_{1}+\bm{\Theta}_{2}+\bm{\Theta}_{3}$ $\mathcal{D}_{1}$$\mathcal{D}_{2}$$\mathcal{D}_{3}$ $\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{2}$ $\frac{1}{2}\bm{N}_{1}-\bm{N}_{2}$ $\bm{\Theta}_{2}-\bm{\Theta}_{3}$ $\bm{N}_{2}-\bm{N}_{3}$ $\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{3}$ $\frac{1}{2}\bm{N}_{1}+\bm{N}_{3}$ $\bm{\Theta}_{1}+\bm{N}_{1}$$\bm{\Theta}_{2}+\bm{N}_{2}$$\bm{\Theta}_{3}+\bm{N}_{3}\endlxSVG@picture$
Figure 1: An overview of the proposed SecCoGC method with simplified notations, when $M=3$, $s=1$.
Each client $k\in\{1,\cdots,K\}\triangleq[K]$ holds a raw local update $\Delta\bm{\Theta}_{k}^{t,I}$ and a secret key $\bm{N}_{k}^{t}$, depicted in Fig. 1.  [^7] In this paper, secret key and privacy noise are used interchangeably. [^8] The secret keys are generated and distributed by a key management server, ensuring that each client has access only to its own keys, without knowledge of their correlations or of any key held by other clients. The secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ are designed to sum to zero,
$\sum_{k=1}^{K}\bm{N}_{k}^{t}=0.$ (6)
To protect data during transmission, each client applies additive masking to its raw local model update,
$\bm{Y}_{k}^{t}=\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t}.$ (7)
Then, each client $k$ transmits $\bm{Y}_{k}^{t}$ to its neighbors $m\in\mathcal{V}_{k}$, where $\mathcal{V}_{k}\triangleq\{m:g_{m,k}\neq 0\}$ (including client $k$ itself), through unreliable link $\tau_{m,k}^{t}$.
In the meanwhile, client $k$ also acts as a relay, listening to its neighboring client $m\in\mathcal{U}_{k}$, where $\mathcal{U}_{k}\triangleq\{m:g_{k,m}\neq 0\}$ (including client $k$ itself). Client $k$ may or may not be able to receive information from client $m$, depending on the realization of link $\tau_{k,m}^{t}$, where
@@
Let $\mathcal{U}_{k}^{t}$ denote the set of updates that client $k$ can successfully learn (including its own updates) at the $t$-th round, i.e., $\mathcal{U}_{k}^{t}=\{m:\tau_{k,m}^{t}=1\}$. Then, client $k$ computes the product $g_{k,m}\bm{Y}_{m}^{t}$ for each $m\in\mathcal{U}_{k}^{t}$, and sums these products to compute the partial sum, i.e.,
@@
If $\mathcal{U}_{k}^{t}=\mathcal{U}_{k}$, client $k$ can successfully decode the message from all of its neighbors, then $\bm{S}_{k}^{t}$ is referred to as a complete partial sum, since the coefficients in the partial sum $\bm{S}_{k}^{t}$ fully match those in $\bm{G}$.
Then, client $k$ sends $\bm{S}_{k}^{t}$ (not necessarily complete) to the server through link $\tau_{k}^{t}$, where
@@
    Combinator: Let $\bm{c}_{f_{t}}$ denote the $f_{t}$-th row in the combination matrix $\bm{C}$. The combinator $\bm{c}_{f_{t}}$ corresponds to $\mathcal{U}^{t}$ should satisfy
@@
    Server Aggregation: If such $\bm{c}_{f_{t}}$ exists, the server can compute the global model by combining these complete partial sums according to $\bm{c}_{f_{t}}$,
@@
It can be verified that the server aggregation in (12) gives exactly the global model pursued by (5) without introducing any additional ambiguity, i.e.,
@@
If the server receives at least KsK-s complete partial sums, it can fully eliminate the privacy noise and accurately reconstruct the global model. Otherwise, clients proceed to the next round using their latest local models by setting $\bm{\Theta}_{k}^{t+1,0}=\bm{\Theta}_{k}^{t,I}$ in (3).
Stopping Rule: The training terminates when the total number of executed rounds reaches $T$, with the final round resulting in a successful reconstruction of the global model. The complete training procedure is detailed in Algorithm 1.
The proposed SecCoGC ensures the correct computation of the global model and reliable cancellation of secret keys in heterogeneous networks through a fixed aggregation structure predetermined by the coding scheme. It not only ensures the optimality of FL algorithms but also protects the user’s privacy at all protocol layers.
Remark III.1 (Encoding and Decoding Complexity).
During client-to-client communication, each client needs to encode its own local updates for transmission and decode $s$ messages from its $s$ neighbors. During client-to-server transmission, at most $M$ clients encode their complete partial sums for transmission, and the server decodes from at most $M$ clients. In total, each client performs $2$ encoding operations and $s$ decoding operations, the server performs at most $M$ decoding operations.
Algorithm 1 Pseudocode of the SecCoGC Protocol
1:Input: $T,K,\eta$, $\bm{\Theta}_{k}^{0}$, $t\leftarrow 0$, $T_{s}$, $R_{t_{s}}\leftarrow 1$, $t_{s}\leftarrow 0$
2:while $\sum_{j=1}^{t_{s}}\leq T$ do
3:  $t\leftarrow t+1$
4:  Client side
5:  for client $k\in[K]$ do
6:   if $\iota=1$ then
7:     $\bm{\Theta}_{k}^{t,0}=\bm{\Theta}^{t-1}$
8:   else
9:     $\bm{\Theta}_{k}^{t+1,0}=\bm{\Theta}_{k}^{t,I}$
10:   end if
11:   for interation $i\in[I]$ do
12:     $\bm{\Theta}_{k}^{t,i}\leftarrow\bm{\Theta}_{k}^{t,i-1}-\eta a_{i}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t,i-1},\bm{\xi}_{k}^{t,i})$
13:   end for
14:   Update $\Delta\bm{\Theta}_{k}^{t,I}=\bm{\Theta}_{k}^{t,I}-\bm{\Theta}_{k}^{t,0}$
15:   Mask $\bm{Y}_{k}^{t}=\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t}$
16:   Send $\bm{Y}_{k}^{t}$ to neighbors in $k\in\mathcal{V}_{k}$
17:   Hear $\bm{Y}_{m}^{t}$ from neighbors in $m\in\mathcal{U}_{k}$
18:   Compute $\bm{S}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{Y}_{m}^{t}$
19:   Send $\bm{S}_{k}^{t}$ to the server
20:  end for
21:  Server side
22:  Detect the combinator $\bm{c}_{f_{t}}$
23:  if $\bm{c}_{f_{t}}$ exists then
24:   $\iota=1$ $\triangleright$ Indicator for GC decoding
25:   $t_{s}\leftarrow t_{s}+1$ $\triangleright$ Index for successful recovery
26:   Update $\bm{\Theta}^{t}=\bm{\Theta}^{t-1}+\frac{1}{K}\sum_{k=1}^{K}c_{f_{t},k}\bm{S}_{k}^{t}$
27:   $R_{t_{s}}\leftarrow 1$
28:  else
29:   $\iota=0$
30:   $R_{t_{s}}\leftarrow R_{t_{s}}+1$ $\triangleright$ Consecutive training rounds
31:  end if
32:  $r\leftarrow r+1$
33:end while
34:Update $T_{s}\leftarrow t_{s}$ $\triangleright$ Total number of global updates
IV Zero-Sum Noise in Real Field
This section begins by discussing the general real-field zero-sum noises. Then, it introduces fairness as a performance metric to ensure equitable privacy protection among clients.
IV-A General Construction
Define $L$ independent randomness components $\bm{Z}_{l}^{t}\sim\mathcal{N}(\mathbf{0},\bm{I}_{D})$,  [^9] The randomness $\bm{Z}_{l}^{t}$ are not necessarily zero-mean Gaussian vectors. It is assumed so for the ease of theoretical analysis. where $l\in[L]$ and $\bm{I}_{D}$ is the identity matrix of size $D\times D$. On each client $k$, the secret key is constructed as a combination of these randomness components, i.e., $\bm{N}_{k}^{t}=\sum_{l=1}^{L}\alpha_{k,l}\bm{Z}_{l}^{t}$. Let $\bm{A}$ (of size $K\times L$) denote the generator matrix comprising all $\alpha_{k,l}$. In matrix form, the secret keys on all clients are generated by
@@ IV-A General Construction
Below, the problem of secure aggregation is reformulated to accommodate the real field.
Condition IV.1 (Real-Field Zero-Sum Secure Aggregation).
Firstly, the secure aggregation scheme must ensure the correct computation of the global model in (6). This implies that each column of the generator matrix $\bm{A}$ must sum to zero, i.e.,
@@
Secondly, any linear combination involving fewer than $K$ (an adjustable threshold) secret keys must be non-trivial, i.e., the zero vector must be excluded from the span of any such subset so that no combination of fewer than $K$ secret keys cancels out to zero. This requires
@@
Thirdly, to avoid zero secret keys, any row in $\bm{A}$ should contain at least one non-zero entry, i.e., $\forall k$, $\exists l:\alpha_{k,l}\neq 0$.
Remark IV.1 (Necessity of the Security Condition in (16)).
Rank K1K-1 is the largest possible rank of $\bm{A}$ while ensuring correctness. It ensures that partial participation does not undermine security, since secret keys cannot be fully canceled. One may argue that the security condition in (16) is unnecessarily strong, considering the fact that GC decoding only leads to either the exact global model reconstruction with full participation or meaningless results. However, considering the client-to-client outage may lead to full rank of the coefficient matrix $\bm{G}$ and allowing for other decoding techniques [35], (16) remains essential under unreliable communication, as it offers privacy guarantees to each local model that hold regardless of the employed decoding strategy.
Lemma IV.1 (Existence Condition).
A generator matrix $\bm{A}\in\mathbb{R}^{K\times L}$ that gives distinct secret keys satisfying Condition IV.1 exists iff
@@
Proof.
The matrix $\bm{A}$ contains KLKL entries, and (15) imposes $L$ linear constraints. Thus, if $KL\geq L$, which holds for all $K$ and $L$, the system has non-zero solutions. Additionally, (16) introduces an extra constraint, requiring $L\geq K-1$. ∎
Construction IV.1 (General Construction).
Condition IV.1 suggests a straightforward construction for $\bm{A}$.
@@
The first K1K-1 secret keys are independently generated with full freedom, while the last key is fully dependent on them. The random generation of the real-field numbers ensures (16) w.p. $1$.
IV-B Fair Zero-Sum Noise Construction
A flexible construction, e.g., Construction IV.1, often leads to variations in privacy noise power. However, signals with low or high magnitudes are susceptible to distortion or clipping by the typically nonlinear power amplifier, resulting in inaccurate and noisy aggregation. In addition, the uneven amount of randomness injected into local models implies unequal levels of privacy preservation across clients. Since the learning weights are predetermined and independent of factors such as communication or privacy, such disparities in noise power are unjustified and may raise fairness concerns.
A concrete example is given to illustrate this issue.
Example IV.1 (Fairness Issue).
Following Construction IV.1 with $K=5$ clients, the last key is generated by $\bm{N}_{5}^{t}=-\sum_{k=1}^{4}\bm{N}_{k}^{t}$, the power of $\bm{N}_{5}^{t}$ can be accumulated significantly. One example can be
@@ IV-B Fair Zero-Sum Noise Construction
The last client receives privacy noise with power 3163-16 times higher than that of the other clients, leading to disproportionately stronger privacy protection despite all clients being assumed equally important.
The condition for fair privacy preservation requires that all privacy noises have identical power  [^10] This paper assumes equal client importance and therefore focuses on uniform privacy preservation. Secret key constructions that provide privacy levels according to heterogeneous learning weights are left for future work., i.e., $\forall k\in[K]:\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$. This leads to the alternative condition below.
Condition IV.2 (Fairness Condition).
@@
Lemma IV.2 (Existence Condition of Fair $\bm{A}$).
A generator matrix $\bm{A}\in\mathbb{R}^{K\times L}$ that gives distinct secret keys and satisfies both the fairness condition (Condition IV.2) and the secure condition (Condition IV.1) iff
@@
Proof.
The fairness condition in (18) imposes $K$ additional constraints. A non-trivial solution for $\bm{A}$ satisfying both (15) and (18) exists if $KL\geq K+L$, i.e., $K\geq 2$, $L\geq 2$. Combining this with Lemma IV.1 completes the proof. ∎
Lemma IV.3 (Lower Bound on Density).
If $\bm{A}$ is of cyclic structure, the number of nonzeros per row should satisfy
@@
To achieve fair zero-sum privacy noise generation, the design must simultaneously satisfy the column-wise linear constraints required for security, and the row-wise $\ell_{2}$-norm constraints required for fairness.
Construction IV.2 (General Fair Construction).
@@
The (21) can be solved by, e.g., the substitution method, the Gröbner Basis Method, etc. However, solving these equations is computationally complex in large systems. Meanwhile, the dual requirements on rows and columns pose a challenge to code construction, as it necessitates the joint design of all rows and columns, rather than treating them separately.
To overcome these limitations, we introduce a simplified generator-matrix construction with a cyclic structure that yields fair and secure secret keys without requiring solving any equations. The core idea is to first set the off-diagonal elements equal, and then determine the diagonal entry so that the $\ell_{2}$-norm across rows is enforced to be equal. In such a design, adjustments to rows do not affect their linear relationships across columns.
Construction IV.3 (Simplified Construction).
 [^11] As a comparison, the covariance-based construction in [16] is tailored to generating privacy noise with a positive-definite correlation matrix using i.i.d. Gaussian randomness. In contrast, the constructions in this paper require neither the positive-definiteness requirement nor the Gaussian assumption.
Let $L=K$ and each row contains $\gamma\geq 1$ (Lemma IV.3) off-diagonal non-zeros. The following construction of $\bm{A}$ yields fair secret keys.
@@
Example IV.2.
Set $\gamma=2$, and $\lambda^{2}=6$, Construction IV.3 gives the following generator matrix,